a model for the path of a toy rocket is given by h=68t-4.9t(t), where h is the altitude in meters and t is the time in seconds. Explain how to find both the maximum altitude of the rocket andhow long it takes to reach that altitude.
Answer by stanbon(75887)

a model for the path of a toy rocket is given by h=68t-4.9t(t), where h is the altitude in meters and t is the time in seconds. Explain how to find both the maximum altitude of the rocket and how long it takes to reach that altitude.
-------------------
There are several ways to find the max of min of a quadratic equation.
---
Max occurs when t = -b/2a
t = -b/2a = -68/(2*-4.9) = 6.93 seconds
---
Max altitude = h(6.93) = 68(6.93)-4.9(6.93^2) = 235.92 meters
--------------------------
Cheers,
Stan H.